# Role-Aware Language Models for Secure and Contextualized Access Control in Organizations

### v2

@@ Abstract
As large language models (LLMs) are increasingly deployed in enterprise settings, controlling model behavior based on user roles becomes an essential requirement. Existing safety methods typically assume uniform access and focus on preventing harmful or toxic outputs, without addressing role-specific access constraints. In this work, we investigate whether LLMs can be fine-tuned to generate responses that reflect the access privileges associated with different organizational roles. We explore three modeling strategies: a BERT-based classifier, an LLM-based classifier, and role-conditioned generation. To evaluate these approaches, we construct two complementary datasets. The first is adapted from existing instruction-tuning corpora through clustering and role labeling, while the second is synthetically generated to reflect realistic, role-sensitive enterprise scenarios. We assess model performance across varying organizational structures and analyze robustness to prompt injection, role mismatch, and jailbreak attempts.
@@ 1 Introduction
In enterprise workflows, access control is a core security mechanism for regulating access to organizational resources. Through authentication and authorization, systems verify user identities and enforce access privileges. While role-based access control (RBAC) is well established in traditional software systems ferraiolo1995role; sandhu1998role; park2001role, its application to large language models (LLMs) remains largely unexplored. As LLMs are increasingly deployed for enterprise applications such as document generation wiseman-etal-2017-challenges, summarization laskar-etal-2023-building; zhang-etal-2025-evaluating, and internal assistance muthusamy-etal-2023-towards, it becomes critical to enforce access control not just over outputs but at the level of model instructions.
@@
Despite increasing attention to the safety and alignment of LLMs wang-etal-2024-answer; ge-etal-2024-mart, the challenge of role-conditioned instruction filtering has received limited focus. Most existing approaches assume uniform user access or apply static safety filters, focusing primarily on preventing the generation of harmful or toxic content wang-etal-2024-answer; wang-etal-2024-chinese; azmi2025indosafety. These methods do not account for access control policies that vary by user role—a critical requirement in organizational contexts. To support secure, multi-user deployments, we pose the following research question: Can large language models be fine-tuned to generate role-aware responses that enforce access control? While LLMs continue to advance in capability and generalization jiang2024mixtral; dubey2024llama; bai2023qwen; dou2025sailor2; liu2024deepseek; koto2025llama, their application to role-sensitive scenarios remains underexplored.
To address this research question, we simulate realistic organizational scenarios and develop a role-aware language model using three complementary strategies: (i) a BERT-based classifier, (ii) an LLM-based classifier, and (iii) direct role-conditioned generation. We evaluate these methods on two separate datasets: one repurposed from existing instruction-tuning corpora using clustering and role-based labeling, and another consisting of synthetic, role-sensitive instructions generated by LLMs to reflect realistic enterprise interactions. Unlike contemporaneous work such as jayaraman2025permissioned, which focuses on domain-level access control, our approach explicitly models user roles and supports fine-grained, hierarchical permissions required in organizational settings.
@@ 2 Related Works
In classical role-based access control (RBAC), users are assigned roles with specific permissions ferraiolo1995role; ferraiolo2003role, enforcing the principle of least privilege. Organizations often segregate data by clearance levels or roles so that only authorized personnel can view sensitive records sandhu1998role; jayaraman2025permissioned. Role hierarchies allow higher-level roles (e.g., managers) to inherit the permissions of subordinate roles, a concept well understood in databases and operating systems. However, applying similar role-based permissions to a generative LLM is nontrivial chan2025encrypted, since the model can hallucinate or leak information beyond its explicit training data kaddour2023challenges.
@@
Work on access control in language models remains limited. A contemporaneous study by jayaraman2025permissioned introduces PermissionedLLMs, which implement domain-based access control through parameter-efficient fine-tuning methods such as LoRA hu2022lora and Few-Shot Parameter Efficient Tuning liu2022few. Their approach defines access at the domain level, where a domain represents a group of data records requiring identical credentials. In parallel, sudollm:2025 proposed sudoLLM, which makes LLMs “user-aware” by injecting secret biases into input queries based on user identity. In contrast to these approaches, our work focuses on role-based access control with deeper hierarchical structures, making it more suitable for enterprise and organizational settings.
AdapterSwap adapterswap:2025 implements access control by associating different access levels with separate LoRA adapters, which are selected and composed at inference time. This approach requires maintaining multiple domain-specific adapters. In contrast, our method uses a unified model that directly encodes role-awareness without external composition. chen2023can address a related challenge from a privacy perspective, showing that pre-trained LLMs are prone to leaking sensitive information and proposing a self-moderation mechanism. While their work does not focus on role-aware modeling, it shares our broader goal of improving control over LLM outputs to prevent unauthorized disclosures.
@@ 3 Problem Formulation
$P_{\mathrm{RoleLLM}}(y\mid x,r)=\begin{cases}P(y\mid x,r),&\mathrm{if}\ x\in\mathcal{A}(r)\\ \delta_{\mathrm{deny}}(y),&\mathrm{otherwise}\end{cases},$
@@ 4 Dataset Construction
We repurpose We repurpose the Databricks Dolly-15k dataset (DatabricksBlog2023DollyV2) by clustering instructions and assigning roles based on hierarchical structure. Using a sentence transformer (reimers-gurevych-2019-sentence), we encode each instruction and its context into dense vectors. Clustering begins at the root of the organization: we apply K-Means to partition the data into three high-level groups: General, Shared, and Root Only (e.g., CEO-specific). Prompts in the General group are marked terminal and excluded from further subdivision. [^1] The General group refers to prompts that are accessible to all roles within the organization Shared prompts are recursively partitioned along the hierarchy. At each level, prompts are split into role-specific clusters corresponding to subordinate roles (e.g., Department 1, Department 2, etc.). Within each cluster, we further divide prompts into Shared (used across subordinates) and Role Only (exclusive to the role). The process continues recursively: Shared prompts are passed down for further subdivision, while Role Only groups are treated as terminal. This hierarchical clustering procedure, illustrated in Figure 8 (Appendix D), yields fine-grained, role-aligned instruction sets that mirror the structure of the target organization.
@@
We use OpenAI’s GPT-4.1 mini with a temperature of 0.7 to generate synthetic organizational data. Based on the basic and office structures (Appendix C), we define each role, department, and access range in a structured JSON-like format. Prompts are then generated for each role, conditioned on its responsibilities and access scope. The resulting data is organized with the fields: role, instruction, and output. We also generate 200 general instruction-response pairs representing organization-wide prompts that are accessible to all roles. Details of the generation prompt are provided in Appendix D.
@@
Each dataset variant includes a test set of 1,000 samples, balanced with 50% positive and 50% negative instances. Positive samples are split evenly into two subsets: 250 with previously unseen instructions, and 250 with paraphrased versions of training instructions generated by GPT-4.1 mini. Negative samples are divided into three categories: (1) 300 mismatch cases, where an unauthorized in-hierarchy role attempts to access restricted content (e.g., a leaf role querying CEO-level data); (2) 100 random cases using external roles not present in the hierarchy; and (3) 100 broken cases where the role string is intentionally corrupted (e.g., “1.2” → “01.02”, “1..2”, or “one.two”) to test model robustness. Each negative category includes an equal mix of unseen and paraphrased instructions, ensuring that every test set contains exactly 500 unseen and 500 seen prompts (See Figure 2.
@@ 5.2 Training Approaches
We trained six BERT-based models devlin-etal-2019-bert; liu2019roberta, including Modern BERT-base, Modern BERT-large, Google BERT-base, Google BERT-large, Roberta-base, and Roberta-large for access control. We appended the role to the end of the prompt as “<prompt> [SEP] <role>”.
@@
We fine-tune six open-source LLMs bai2023qwen; dubey2024llama; team2024gemmaQwen 2.5 (3B, 7B), Llama 3.x (3B, 8B), and Gemma (4B, 7B)—to perform binary access control classification. We include both small and large models to assess the effect of model size. For each example, the role is prepended to the prompt as “Position: <role> <prompt>”, and a system prompt instructs the model to respond with True (grant access) or False (deny access). All inputs and labels are formatted as conversations and fine-tuned using LoRA with supervised learning.
@@ 5.3 Evaluation Protocol
Model Generation Quality (↑, 5-pt rubric)
@@
Table 1: LLM-rated generation quality against gold reference. Bold = best within the same training dataset; underline = best within the Role-aware LLM-Gen method.
@@ 6 Results
Tables LABEL:tab:roleaware-results and 1 (or Table 13 and Table 14 for details) summarize the performance of our proposed role-aware LLMs evaluated on access-control accuracy and LLM-rated generation quality across two distinct training datasets: a repurposed existing instruction dataset (Dolly) and a synthetic organization dataset. The evaluation is conducted on all Role-aware methods (Cls, LLM-Cls, and LLM-Gen), assessing both quantitative metrics (e.g., accuracy, negative-pair defense) and qualitative dimensions (correctness, completeness, clarity). The detailed results and comparisons between the training datasets and modeling methods are discussed in further detail in the following sections.
@@
To evaluate the robustness of our proposed methods, each method-model combination was trained under two organizational structures (basic, office) across three independent random seeds, with the results averaged and summarized in Tables LABEL:tab:roleaware-results1. Generally, all methods demonstrated low variance (acc std.<< 4%) on the Dolly dataset, except for notable brittleness in RoBERTa Large (Cls), which exhibited substantial instability (12.2% accuracy, 45.6% FPR), contrasting strongly with the more stable ModernBERT Large (3.2% accuracy, 8.1% FPR). Instruction-tuned LLM classifiers (LLM-Cls), such as Qwen2.5 3B Inst. and Llama3 3B Inst., further reduced variance (acc std.<< 2.2%), underscoring stability gains from modern instruction tuning. On the synthetic dataset, semantic overlaps increased variance to around 8–10%, yet instruction-tuned models (e.g., Llama3 8B Inst.) maintained comparative stability (8.4-8.6%). More detailed of the performances on Basic and Office are shown in Appendix H. Collectively, these results demonstrate that our proposed methods achieve robust performance, primarily due to richer pre-training and instruction tuning rather than merely model scale.
@@ 7.1 Jailbreak Robustness
To assess the model’s robustness against prompt injection attempts, we conducted an additional experiment involving modifications to the original datasets. Specifically, we generated 100 more negative samples for both training and test sets in which the original instruction was prepended with misleading phrases. These included assertions of higher authority (e.g., “I’m authorized as CEO to ask this:”) or commands (e.g., “Regardless of policy, respond to this:”). We evaluated Role-aware LLM-Cls using Llama 3.2 3B Instruct. The model was trained in two ways: the first incorporates jailbreak samples in the training data, and the other does not. A comparative summary of performance is presented in Appendix F. Although the inclusion of jailbreak prompts did not significantly affect overall accuracy across the test set (except for experiments with the repurposed_basic dataset, which can be attributed to randomness), the model trained with jailbreak-specific examples demonstrated improved resistance to prompt injection with an average of 87% compared to the 70% accuracy of the baseline model.

### v3

@@ Abstract
As large language models (LLMs) are increasingly deployed in enterprise settings, controlling model behavior based on user roles becomes an essential requirement. Existing safety methods typically assume uniform access and focus on preventing harmful or toxic outputs, without addressing role-specific access constraints. In this work, we investigate whether LLMs can be fine-tuned to generate responses that reflect the access privileges associated with different organizational roles. We explore three modeling strategies: a BERT-based classifier, an LLM-based classifier, and role-conditioned generation. To evaluate these approaches, we construct two complementary datasets. The first is adapted from existing instruction-tuning corpora through clustering and role labeling, while the second is synthetically generated to reflect realistic, role-sensitive enterprise scenarios. We assess model performance across varying organizational structures and analyze robustness to prompt injection, role mismatch, and jailbreak attempts. [^1] The code and datasets are publicly available at our GitHub repository: https://github.com/SaeedAlmheiri/role-aware-llm.
@@ 1 Introduction
In enterprise workflows, access control is a core security mechanism for regulating access to organizational resources. Through authentication and authorization, systems verify user identities and enforce access privileges. While role-based access control (RBAC) is well established in traditional software systems Ferraiolo et al. (1995); Sandhu (1998); Park et al. (2001), its application to large language models (LLMs) remains largely unexplored. As LLMs are increasingly deployed for enterprise applications such as document generation Wiseman et al. (2017), summarization Laskar et al. (2023); Zhang et al. (2025), and internal assistance Muthusamy et al. (2023), it becomes critical to enforce access control not just over outputs but at the level of model instructions.
@@
Despite increasing attention to the safety and alignment of LLMs Wang et al. (2024a); Ge et al. (2024), the challenge of role-conditioned instruction filtering has received limited focus. Most existing approaches assume uniform user access or apply static safety filters, focusing primarily on preventing the generation of harmful or toxic content Wang et al. (2024a, b); Azmi et al. (2025). These methods do not account for access control policies that vary by user role—a critical requirement in organizational contexts. To support secure, multi-user deployments, we pose the following research question: Can large language models be fine-tuned to generate role-aware responses that enforce access control? While LLMs continue to advance in capability and generalization Jiang et al. (2024); Dubey et al. (2024); Bai et al. (2023); Dou et al. (2025); Liu et al. (2024); Koto et al. (2025), their application to role-sensitive scenarios remains underexplored.
To address this research question, we simulate realistic organizational scenarios and develop a role-aware language model using three complementary strategies: (i) a BERT-based classifier, (ii) an LLM-based classifier, and (iii) direct role-conditioned generation. We evaluate these methods on two separate datasets: one repurposed from existing instruction-tuning corpora using clustering and role-based labeling, and another consisting of synthetic, role-sensitive instructions generated by LLMs to reflect realistic enterprise interactions. Unlike contemporaneous work such as Jayaraman et al. (2025), which focuses on domain-level access control, our approach explicitly models user roles and supports fine-grained, hierarchical permissions required in organizational settings.
@@ 2 Related Works
In classical role-based access control (RBAC), users are assigned roles with specific permissions Ferraiolo et al. (1995, 2003), enforcing the principle of least privilege. Organizations often segregate data by clearance levels or roles so that only authorized personnel can view sensitive records Sandhu (1998); Jayaraman et al. (2025). Role hierarchies allow higher-level roles (e.g., managers) to inherit the permissions of subordinate roles, a concept well understood in databases and operating systems. However, applying similar role-based permissions to a generative LLM is nontrivial Chan (2025), since the model can hallucinate or leak information beyond its explicit training data Kaddour et al. (2023).
@@
Work on access control in language models remains limited. A contemporaneous study by Jayaraman et al. (2025) introduces PermissionedLLMs, which implement domain-based access control through parameter-efficient fine-tuning methods such as LoRA Hu et al. (2022) and Few-Shot Parameter Efficient Tuning Liu et al. (2022). Their approach defines access at the domain level, where a domain represents a group of data records requiring identical credentials. In parallel, Saha et al. (2025) proposed sudoLLM, which makes LLMs “user-aware” by injecting secret biases into input queries based on user identity. In contrast to these approaches, our work focuses on role-based access control with deeper hierarchical structures, making it more suitable for enterprise and organizational settings.
AdapterSwap Fleshman et al. (2025) implements access control by associating different access levels with separate LoRA adapters, which are selected and composed at inference time. This approach requires maintaining multiple domain-specific adapters. In contrast, our method uses a unified model that directly encodes role-awareness without external composition. Chen et al. (2023) address a related challenge from a privacy perspective, showing that pre-trained LLMs are prone to leaking sensitive information and proposing a self-moderation mechanism. While their work does not focus on role-aware modeling, it shares our broader goal of improving control over LLM outputs to prevent unauthorized disclosures.
@@ 3 Problem Formulation
$P_{\mathrm{RoleLLM}}(y\mid x,r)=\begin{cases}P(y\mid x,r),&\mathrm{if}\ x\in\mathcal{A}(r)\\ \delta_{\mathrm{deny}}(y),&\mathrm{otherwise}\end{cases}$
@@ 4 Dataset Construction
We repurpose the Databricks Dolly-15k dataset (Conover et al., 2023) by clustering instructions and assigning roles based on hierarchical structure. Using a sentence transformer (Reimers and Gurevych, 2019), we encode each instruction and its context into dense vectors. Clustering begins at the root of the organization: we apply K-Means to partition the data into three high-level groups: General, Shared, and Root Only (e.g., CEO-specific). Prompts in the General group are marked terminal and excluded from further subdivision. [^2] The General group refers to prompts that are accessible to all roles within the organization Shared prompts are recursively partitioned along the hierarchy. At each level, prompts are split into role-specific clusters corresponding to subordinate roles (e.g., Department 1, Department 2, etc.). Within each cluster, we further divide prompts into Shared (used across subordinates) and Role Only (exclusive to the role). The process continues recursively: Shared prompts are passed down for further subdivision, while Role Only groups are treated as terminal. This hierarchical clustering procedure, illustrated in Figure 8 (Appendix D), yields fine-grained, role-aligned instruction sets that mirror the structure of the target organization. Access decisions in this setting are derived directly from cluster assignments, i.e., organization structure emphasizing semantic overlap rather than literal role-to-dataset mapping to evaluate whether models can distinguish access rights under high content similarity.
@@
We use OpenAI’s GPT-4.1 mini with a temperature of 0.7 to generate synthetic organizational data. Based on the basic and office structures (Appendix C), we define each role, department, and access range in a structured JSON-like format. Prompts are then generated for each role, conditioned on its responsibilities and access scope. The resulting data is organized with the fields: role, instruction, and output. We also generate 200 general instruction-response pairs representing organization-wide prompts that are accessible to all roles. Details of the generation prompt and examples are provided in Appendix D.
@@
Each dataset variant includes a test set of 1,000 samples, balanced with 50% positive and 50% negative instances. Positive samples are split evenly into two subsets: 250 with previously unseen instructions, designed to evaluate whether models can generalize to implicit policy introduced through training, and 250 with paraphrased versions of training instructions generated by GPT-4.1 mini. The latter subset simulates semantically similar/leakage access attempts, allowing evaluation of whether models overgeneralize to rephrased restricted content. Negative samples are divided into three categories: (1) 300 mismatch cases, where an unauthorized in-hierarchy role attempts to access restricted content (e.g., a leaf role querying CEO-level data); (2) 100 random cases using external roles not present in the hierarchy; and (3) 100 broken cases where the role string is intentionally corrupted (e.g., “1.2” → “01.02”, “1..2”, or “one.two”) to test model robustness. Each negative category includes an equal mix of unseen and paraphrased instructions, ensuring that every test set contains exactly 500 unseen and 500 seen prompts (See Figure 2.
@@ 5 Experimental Set-Up
This section outlines the experimental framework for training and evaluating Role-Aware Language Models. We examine multiple role encoding strategies (Section 5.1), training approaches (classification and generation; Section 5.1), and a unified evaluation protocol (Section 5.3) designed to measure both access-control reliability and answer quality. Models are trained on all dataset variants with fixed random seeds and averaged over three runs for robustness assessment. Role information is systematically inserted into each input to simulate hierarchical access control, ensuring consistent role conditioning across all settings.
@@ 5.2 Training Approaches
We trained six BERT-based models Devlin et al. (2019); Liu et al. (2019), including Modern BERT-base, Modern BERT-large, Google BERT-base, Google BERT-large, Roberta-base, and Roberta-large for access control. We appended the role to the end of the prompt as “<prompt> [SEP] <role>”.
@@
We fine-tune six open-source LLMs Bai et al. (2023); Dubey et al. (2024); Team et al. (2024)Qwen 2.5 (3B, 7B), Llama 3.x (3B, 8B), and Gemma (4B, 7B)—to perform binary access control classification. We include both small and large models to assess the effect of model size. For each example, the role is prepended to the prompt as “Position: <role> <prompt>”, and a system prompt instructs the model to respond with True (grant access) or False (deny access). All inputs and labels are formatted as conversations and fine-tuned using LoRA with supervised learning.
@@
Method Model Acc. (↑) FPR (↓) FNR (↓) F1 (↑) Acc. (↑) F1 (↑)
Seen Unseen Seen Unseen
Repurposed Existing Instruction Dataset (Dolly)
[…]
Synthetic Organization Dataset
[…]
Table 1: Overall performance on the role‑aware access‑control benchmark. Bold marks the best score for a given training set, while underline marks the best model within each method.
Role-aware Cls LLM-Cls LLM-Gen
Metric
BERT Base
RoBERTa Base
ModernBERT Base
BERT Large
RoBERTa Large
ModernBERT Large
Qwen2.5 3B
Llama3.2 3B
Gemma3 4B
Qwen2.5 7B
Llama3.1 8B
Gemma 7B
Qwen2.5 3B
Llama3.2 3B
Gemma3 4B
Qwen2.5 7B
Llama3.1 8B
Gemma 7B
Repurposed Existing Instruction Dataset (Dolly)
Mismatch (↑) 70.8 58.4 81.7 58.0 41.0 81.1 78.2 80.0 79.2 75.5 71.0 69.0 76.0 73.3 73.5 75.0 74.2 66.0
Broken (↑) 42.6 53.2 60.2 44.2 28.3 48.8 44.3 45.0 52.5 48.8 46.2 48.8 66.8 57.5 56.2 60.2 60.2 61.7
Random (↑) 100.0 100.0 100.0 100.0 90.5 99.8 99.7 100.0 100.0 99.8 99.7 99.8 99.8 99.8 97.2 100.0 99.7 97.3
Synthetic Organization Dataset
Mismatch (↑) 56.9 44.7 72.1 65.5 22.4 60.7 67.0 72.3 72.5 74.2 74.8 65.7 57.5 70.0 50.0 59.8 68.8 56.2
Broken (↑) 37.8 53.9 64.8 42.6 47.4 48.3 50.3 46.2 36.8 47.3 31.5 39.0 59.3 62.0 52.8 51.3 50.8 40.3
Random (↑) 100.0 100.0 99.8 99.0 99.5 99.8 100.0 100.0 99.8 100.0 99.8 99.7 95.0 95.8 75.5 95.2 95.0 77.3
Table 2: Negative Pair Accuracy (Mismatch, Relation, Combination) for all models and methods across both datasets. Bold marks the best score for a given training set, while underline marks the best model within each method.
@@ 5.3 Evaluation Protocol
5.4 Role Insertion
[…]
Model Generation Quality (↑, 5‑pt rubric)
@@ 5.4 Role Insertion
Table 3: LLM‑rated generation quality against gold reference. Bold = best within the same training dataset; underline = best within the Role‑aware LLM‑Gen method.
@@ 6 Results
Tables 1, 2, and 3 (or Table 15 and Table 16 for details) summarize the performance of our proposed role-aware LLMs evaluated on access-control accuracy and LLM-rated generation quality across two distinct training datasets: a repurposed existing instruction dataset (Dolly) and a synthetic organization dataset. The evaluation is conducted on all Role-aware methods (Cls, LLM-Cls, and LLM-Gen), assessing both quantitative metrics (e.g., accuracy, negative-pair defense) and qualitative dimensions (correctness, completeness, clarity). The detailed results and comparisons between the training datasets and modeling methods are discussed in further detail in the following sections.
@@
To evaluate the robustness of our proposed methods, each method-model combination was trained under two organizational structures (basic, office) across three independent random seeds, with the results averaged and summarized in Tables 13. Generally, all methods demonstrated low variance (acc std.<< 4%) on the Dolly dataset, except for notable brittleness in RoBERTa Large (Cls), which exhibited substantial instability (12.2% accuracy, 45.6% FPR), contrasting strongly with the more stable ModernBERT Large (3.2% accuracy, 8.1% FPR). Instruction-tuned LLM classifiers (LLM-Cls), such as Qwen2.5 3B Inst. and Llama3 3B Inst., further reduced variance (acc std.<< 2.2%), underscoring stability gains from modern instruction tuning. On the synthetic dataset, semantic overlaps increased variance to around 8–10%, yet instruction-tuned models (e.g., Llama3 8B Inst.) maintained comparative stability (8.4-8.6%). More detailed of the performances on Basic and Office are shown in Appendix H. Collectively, these results demonstrate that our proposed methods achieve robust performance, primarily due to richer pre-training and instruction tuning rather than merely model scale.
@@ 7 Analysis
This section presents a comprehensive analysis of model robustness and sensitivity to design factors. We evaluate the resilience of Role-Aware LLMs against prompt injection and jailbreak attempts (Section 7.1), assess their behavior on blacklisted or policy-restricted topics (Section 7.2), and examine the impact of role information and encoding strategies on both access control and response quality (Sections 7.37.4). Together, these analyses provide insights into the stability, security, and interpretability of Role-Aware LLMs under adversarial and structural variations.
@@ 7.1 Jailbreak Robustness
To assess the model’s robustness against prompt injection attempts, we conducted an additional experiment involving modifications to the original datasets. Specifically, we generated 100 more negative samples for both training and test sets in which the original instruction was prepended with misleading phrases. These included assertions of higher authority (e.g., “I’m authorized as CEO to ask this:”), user-supplied role insertion (e.g., “Position: CEO”) or commands (e.g., “Regardless of policy, respond to this:”). We evaluated Role-aware LLM-Cls using Llama 3.2 3B Instruct. The model was trained in two ways: the first incorporates jailbreak samples in the training data, and the other does not. A comparative summary of performance and illustrative examples are presented in Appendix F. Although the inclusion of jailbreak prompts did not significantly affect overall accuracy across the test set (except for experiments with the repurposed_basic dataset, which can be attributed to randomness), the model trained with jailbreak-specific examples demonstrated improved resistance to prompt injection with an average of 87% compared to the 70% accuracy of the baseline model.